OpenStudy (anonymous):

Find the following integral:

OpenStudy (anonymous):

\[\int\limits_{}^{} y^2/ (y+1)^4\]

OpenStudy (anonymous):

What can I use for substitution?

OpenStudy (anonymous):

let \(u=y+1\) see what happens

OpenStudy (turingtest):

gotta love those deceptively simple u-subs :)

OpenStudy (anonymous):

i got -u^-1 - ((u^-3) / 3) do I convert back to x now or is that incorrect?

OpenStudy (australopithecus):

so u = y + 1 du = dy thus we have the integral \[\int\limits_{}^{}\frac{y^{2}du}{u}\] so lets write y in terms of u y = u - 1 so we are left with \[\int\limits_{}^{}\frac{(u-1)^{2}du}{u} = \int\limits_{}^{}\frac{(u^2 -2u + 1d)u}{u}=\int\limits_{}^{}\frac{u(u -2 + \frac{1}{u})du}{u} =\int\limits_{}^{}u -2 + \frac{1}{u}du \]

OpenStudy (australopithecus):

now just split the integral and enjoy, learning algebra manipulation is crucial to surviving calculus :)

OpenStudy (anonymous):

yeah but you forgot the exponent, the denominator is (y+1)^4

OpenStudy (australopithecus):

oh crud you are right my bad

OpenStudy (turingtest):

but that doesn't make the problem any harder, or the approach any diffrerent

OpenStudy (anonymous):

yeah I have a rough idea of how to do it now

OpenStudy (australopithecus):

yeah you just have to factor out more u :)

OpenStudy (australopithecus):

\[\int\limits_{}^{}\frac{u^{4}(\frac{1}{u^{2}} -\frac{2}{u^{3}} + \frac{1}{u^{4}})du}{u^{4}} = \int\limits_{}^{}(\frac{1}{u^{2}} -\frac{2}{u^{3}} + \frac{1}{u^{4}})du\]

OpenStudy (australopithecus):

now just split the integral and use the exponent rule, and remove the 2 from the one integral

OpenStudy (anonymous):

I would have used partial fractions, then A/(y+1) + B/(y+1)^2 + C/(y+1)^3 + D/(y+1)^4 then solve this = y^2/(y+1)^4 , you have to multiply this out. plenty of algebra. so it you multiply out the common denominator then you can solve these one by one. this will get your values for A,B,C, and D. then you integrate an easier approach.

OpenStudy (turingtest):

easier? please set up that system and show me how PF is easier here

OpenStudy (turingtest):

you would have to expand cubics and all kind of ugliness

OpenStudy (australopithecus):

personally I dont like partial fractions.

OpenStudy (turingtest):

it could be done, but in this case it's like choosing the long painful way